OpenStudy (anonymous):

a+3/a^2+a-12 (divided by) a^2-9/a^2+7a+12

OpenStudy (anonymous):

\[ \ \ \ \ \frac{a+3}{a^2+a-12}\div\frac{a^2-9}{a^2+7a+12}\\ =\frac{a+3}{(a+4)(a-3)}\times\frac{(a+4)(a+3)}{(a+3)(a-3)}\\ =\frac1{a-3}\times\frac{a+3}{a-3}\\ =\frac{a+3}{a^2-6a+9} \]

OpenStudy (anonymous):

omg thanku :-))) @oldrin.bataku